OpenStudy (anonymous):

Lucy is k years old. She has a sister who is 3 yea younger than she is and another sister who is five years less than twice Lucy's age. Write an expression, in simplififed form, for the sum of the three girls' ages. Would 4k-8 be right?

OpenStudy (anonymous):

ages1 = k ages2 = k-3 ages3 = 2k - 5 ages1 +ages2 +ages3 = k + k - 3 + 2k -5 = 4k - 8

OpenStudy (anonymous):

Thanks for the help!!